Chemistry




How much heat is absorbed in the complete reaction of 3.00 grams of SiO2 with excess carbon in the reaction below? DH0 for the reaction is +624.6 kJ.

SiO2(g) + 3C(s) ® SiC(s) + 2CO(g)






A) 366 kJ

B.) 1.13 ´ 105 kJ

C.) 5.06 kJ

D.) 1.33 ´ 104 kJ

E.) 31.2 kJ
Answer #1
60.08 g is mm of sio2
moles of SiO2=3/60.08=.05 mole
heat absorbed=624.6*.05
=31.24 kj
Answer #2
31.2 kJ
answered by: LASTDAYS
Answer #3
B.) 1.13 ´ 105 kJ
Answer #4
Molar mass of SiO2 = 28 + 2*16 = 60 g / mol
SiO2(g) + 3C(s) --> SiC(s) + 2CO(g) ; dHo = +624.6 KJ

According to the equation ,
60 g of SiO2 requires 624.6KJ of energy
3 g of SiO2 requires X KJof energy
X = ( 624.6 * 3 ) / 60
=
31.23 KJ of energy
answered by: Anonymous 2